### v1

# The lattice structure of negative Sobolev and extrapolation spaces

Abstract.
It is well-known that the Sobolev spaces $W^{k,p}(\mathbb{R}^{d})$ are vector lattices with respect to the pointwise almost everywhere order if $k\in\{0,1\}$, but not if $k\geq 2$. In this note, we consider negative $k$ and show that the span of the positive cone in $W^{k,p}(\mathbb{R}^{d})$ is a vector lattice in this case.
We also prove a related abstract result: if $(T(t))_{t\in[0,\infty)}$ is a positive $C_{0}$-semigroup on a Banach lattice $X$ with order continuous norm, then the span of the cone $X_{-1,+}$ in the extrapolation space $X_{-1}$ is a vector lattice. This complements results obtained by Bátkai, Jacob, Wintermayr, and Voigt in the context of perturbation theory and provides additional context for the theory of infinite-dimensional positive systems.
Key words and phrases:
ordered Banach spaces; Sobolev spaces; extrapolation space; Banach lattice; vector lattice
2020 Mathematics Subject Classification:
46B40; 46B42; 46E35
1. Introduction
For every $p\in[1,\infty]$, the first order Sobolev space $W^{1,p}(\mathbb{R}^{d})$ is a vector lattice with respect to the pointwise almost everywhere order. This follows from the fact that $\left\lvert f\right\rvert\in W^{1,p}(\mathbb{R}^{d})$ for all $f\in W^{1,p}(\mathbb{R}^{d})$ [9, Lemma 7.6]; in other words, the Sobolev space is even a sublattice of the Banach lattice $L^{p}(\mathbb{R}^{d})$. In contrast, $W^{k,p}(\mathbb{R}^{d})$ is not a vector lattice for $k\geq 2$ [2, Example 2.3(d)]; however, the cone is generating. In this note we are interested in the case of negative $k$. At first glance, the situation seems to be trivial: for an integer $k<0$, there are also elements of $W^{k,p}(\mathbb{R}^{d})$ which cannot be written as the difference of two positive elements. In particular, the cone is not generating, whence the space is not a vector lattice.
However this simple answer does not really give us much insight into the order structure of $W^{k,p}(\mathbb{R}^{d})$ for negative $k$ – it simply means that the cone is small compared to the entire space (which is, from a theoretical point of view, a consequence of the fact that the cone in Sobolev spaces of positive order is not normal). If we restrict our attention to the span $W^{k,p}(\mathbb{R}^{d})_{+}-W^{k,p}(\mathbb{R}^{d})_{+}$ of the positive cone, we shall see in Section 3 that this is indeed a vector lattice for every integer $k<0$.
A related phenomenon occurs, in a more abstract setting, in the theory of positive operator semigroups. If $X$ is a Banach lattice and $T=(T(t))_{t\geq 0}$ is a positive $C_{0}$-semigroup on $X$, one can associate a so-called extrapolation space $X_{-1}$ to $T$ – this is a larger space into which $X$ embeds continuously and densely. The space $X_{-1}$ can be considered as an abstract Sobolev space of negative order. To get an order structure on $X_{-1}$ it is natural to define the cone $X_{-1,+}$ as the closure of $X_{+}$ within the space $X_{-1}$. This order structure was introduced by Bátkai, Jacob, Wintermayr, and Voigt in [4] in order to study positive unbounded perturbations of semigroup generators; they also proved that this is indeed a cone and that $X_{+}=X_{-1,+}\cap X$, i.e., the order structures on $X$ and $X_{-1}$ are compatible [4, Proposition 2.3]. The same order on $X_{-1}$ plays a significant role in the theory of positive linear systems in infinite-dimensions: there, one is interested in positive control operators $B$ that operate from an ordered Banach space $U$, the so-called input space, into the extrapolation space $X_{-1}$; see the recent papers [7, 3] and also [8]. Just as for the concrete Sobolev space $W^{k,p}(\mathbb{R}^{d})$ for $k<0$, the cone $X_{-1,+}$ does not span $X_{-1}$, in general, and hence $X_{-1}$ is not a vector lattice. This was observed in a concrete example in [4, Example 5.1]. Yet, we show in Section 4 that a similar phenomenon as for the concrete Sobolev spaces occurs, i.e., the span $X_{-1,+}-X_{-1,+}$ is indeed a vector lattice provided that the Banach lattice $X$ has order continuous norm.
Our results are based on general results on how to transfer lattice properties between ordered Banach spaces that are not isomorphic. We prove those results in Section 2. Different results regarding when an ordered Banach space (or ordered vector space) is automatically a vector lattice can be found in [10, Section 4].
We close this introduction by recalling of few notions from the theory of ordered Banach spaces and Banach lattices.
Preliminaries on ordered Banach spaces and Banach lattices
Let $X$ be a real vector space. A non-empty set $X_{+}\subseteq X$ is called a cone if $\alpha X_{+}+\beta X_{+}\subseteq X_{+}$ for all real numbers $\alpha,\beta\geq 0$ and $X_{+}\cap(-X_{+})=\{0\}$. The real vector space $X$ together with a cone $X_{+}$ is called an ordered vector space and $X_{+}$ is called the positive cone of $X$. The term ordered space is justified by the fact that $X_{+}$ induces a partial order on $X$ that is compatible with the vector space structure: one sets $x\leq y$ for $x,y\in X$ if and only if $y-x\in X_{+}$. So in particular, $x\geq 0$ if and only if $x\in X_{+}$. For this reason, the elements of $X_{+}$ are called the positive elements of $X$. We call the cone of $X$ generating if $X=X_{+}-X_{+}$. For $x,z\in X$, the order interval $[x,z]$ is defined as the set $\{y\in X:x\leq y\leq z\}$; note that $[x,z]$ is non-empty if and only if $x\leq z$. A non-empty subset $C\subseteq X_{+}$ is called a face of $X_{+}$ if it is also a cone and if $0\leq x\leq y$ in $X$ and $y\in C$ implies $x\in C$.
By an ordered Banach space, we mean a Banach space $X$ which is also an ordered vector space whose positive cone $X_{+}$ is closed with respect to the norm topology. The cone of an ordered Banach space is called normal if there exists a number $M\geq 1$ such that $0\leq x\leq y$ implies $\left\lVert x\right\rVert\leq M\left\lVert y\right\rVert$ for all $x,y\in X$. If $X$ is an ordered Banach space, then it is easy to see that $\operatorname{span}X_{+}=X_{+}-X_{+}$. Endowing this space with the norm
$\left\lVert x\right\rVert_{\operatorname{span}X_{+}}:=\inf\{\left\lVert y\right\rVert+\left\lVert z\right\rVert:\,y,z\in X_{+}\text{ and }x=y-z\}$ (1.1)
makes $\operatorname{span}X_{+}$ an ordered Banach space and makes the embedding $\operatorname{span}X_{+}\hookrightarrow X$ continuous [2, Lemma 2.2]. Moreover, on the cone $X_{+}$ this norm coincides with the norm on $X$. Hence, if the cone $X_{+}$ is normal in $X$, then it is also normal with respect to $\left\lVert\mathord{\,\cdot\,}\right\rVert_{\operatorname{span}X_{+}}$.
An ordered vector space $X$ is called a vector lattice if any two elements have a supremum (equivalently, an infimum). The positive cone of a vector lattice is always generating and each element $x\in X$ has a modulus that is defined as $\left\lvert x\right\rvert:=\sup\{-x,x\}$. An ordered Banach space is called a Banach lattice if it is a vector lattice and the order is compatible with the norm in the following sense that
$\left\lvert x\right\rvert\leq\left\lvert y\right\rvert\Rightarrow\left\lVert x\right\rVert\leq\left\lVert y\right\rVert$
for all $x,y\in X$. The cone of a Banach lattice is always normal. If $X$ is an ordered Banach space and $\operatorname{span}X_{+}$ is norm dense in $X$, then the so called dual wedge
$X^{\prime}_{+}:=\{x^{\prime}\in X^{\prime}:\,\langle x^{\prime},x\rangle\geq 0\text{ for all }x\in X_{+}\}.$
is also a cone and thus turns the dual space $X^{\prime}$ into an ordered Banach space (if $\operatorname{span}X_{+}$ is not dense in $X$, then $X^{\prime}_{+}\cap-X^{\prime}_{+}$ will may not be zero, so $X^{\prime}$ is then only a pre-ordered Banach space). If $X$ is a Banach lattice, then the dual space is a Banach lattice as well.
For ordered Banach spaces $X$ and $Z$ we denote the space of all bounded linear operators from $X$ to $Z$ by $\mathcal{L}(X,Z)$ and we abbreviate $\mathcal{L}(X):=\mathcal{L}(X,Z)$. A linear map $T:X\to Z$ is called positive if $TX_{+}\subseteq Z_{+}$. As a consequence of the Hahn-Banach separation theorem, this is equivalent to $T^{\prime}Z^{\prime}_{+}\subseteq X^{\prime}_{+}$. By $\mathcal{L}(X,Z)_{+}$ we denote the set of all positive operators in $\mathcal{L}(X,Z)$. The identity operator on $X$ will be denoted by $\operatorname{id}_{X}$.
2. Abstract results on ordered Banach spaces
Recall that a Banach lattice $Z$ is called monotonically complete if every increasing norm-bounded net in $Z$ (equivalently, in $Z_{+}$) has a supremum [11, Definition 2.4.18]. Every $L^{p}$-space is monotonically complete for $p\in[1,\infty)$ and $L^{\infty}$-spaces over a $\sigma$-finite measure space are also monotonically complete. Clearly, a monotonically complete Banach lattice is Dedekind complete, i.e., every order-bounded increasing net in such a space has a supremum. The space $c_{0}$ of sequences that converge to 00 shows that the converse implication does not hold. A KB-space is a Banach lattice in which every increasing norm-bounded net (equivalently, sequence) is norm-convergent. Typical examples of KB-spaces are all reflexive Banach lattices and $L^{1}$-spaces (i.e., all $L^{p}$-spaces for $1\leq p<\infty$).
Theorem 2.1.
Let $X$ be an ordered Banach space with generating cone and let $Z$ be a Banach lattice. Assume that there exists an operator $J\in\mathcal{L}(X,Z)_{+}$ and a sequence $(R_{n})\subseteq\mathcal{L}(Z,X)_{+}$ such that $(R_{n}J)\subseteq\mathcal{L}(X)$ converges to the $\operatorname{id}_{X}$ with respect to the weak operator topology.
Then $\operatorname{span}(X^{\prime}_{+})=X^{\prime}_{+}-X^{\prime}_{+}$ is a monotonically complete Banach lattice with respect to a norm equivalent $\left\lVert\mathord{\,\cdot\,}\right\rVert_{\operatorname{span}X^{\prime}_{+}}$. In addition, if $X$ is reflexive then the Banach lattice $\operatorname{span}(X^{\prime}_{+})$ is a KB-space.
We point out that, in case of $X$ being reflexive, the conclusion that $\operatorname{span}X^{\prime}_{+}$ is a KB-space cannot be strengthened to $\operatorname{span}X^{\prime}_{+}$ also being reflexive; see the discussion after Corollary 2.4 below. Note that, if a sequence in $X_{+}$ converges in $X$, then it may not necessarily converge in $\operatorname{span}X_{+}$ as well. However, this is the case for increasing sequences. We state this observation in the following lemma as it is used multiple times in the sequel.
Lemma 2.2.
Let $X$ be an ordered Banach space and let $(x_{n})$ be an increasing net of elements in $\operatorname{span}X$. If $(x_{n})$ is norm-convergent in $X$, then it is also norm convergent with respect to $\left\lVert\mathord{\,\cdot\,}\right\rVert_{\operatorname{span}X_{+}}$.
Proof.
Let $(x_{n})$ be norm convergent in $X$, to say $x$. Since $(x_{n})$ is increasing, $x-x_{n}\in X_{+}$ for each index $n$. As the norm on $X$ and the norm on $\operatorname{span}X_{+}$ coincide on $X_{+}$, we obtain that
$\left\lVert x_{n}-x\right\rVert_{\operatorname{span}X_{+}}=\left\lVert x_{n}-x\right\rVert_{X}\to 0$
as $n\to\infty$. ∎
Proof of Theorem 2.1.
Step 1: The span is a lattice. Let $x^{\prime}\in\operatorname{span}X^{\prime}_{+}$. It suffices to prove that ±$\pm x^{\prime}$ has a supremum in $\operatorname{span}X^{\prime}_{+}$. Since $Z$ is a Banach lattice, so is its dual space $Z^{\prime}$. Hence, we can take the modulus $\left\lvert R_{n}^{\prime}x^{\prime}\right\rvert$ in $Z^{\prime}$ for each index $n$. We show that the sequence $\big{(}J^{\prime}\left\lvert R_{n}^{\prime}x^{\prime}\right\rvert\big{)}$ is norm-bounded in $X^{\prime}$. First of all, because $(R_{n}J)$ converges weakly to $\operatorname{id}_{X}$, the sequence $(J^{\prime}R_{n}^{\prime})$ converges to $\operatorname{id}_{X^{\prime}}$ in the weak-operator topology. In particular, it is bounded. Secondly, as $x^{\prime}\in X^{\prime}_{+}-X^{\prime}_{+}$, there exists a vector $y\in X^{\prime}_{+}$ such that ±$\pm x^{\prime}\leq y^{\prime}$. Therefore, ±$\pm R_{n}^{\prime}x^{\prime}\leq R_{n}^{\prime}y^{\prime}$ in $Z^{\prime}$ and hence, $J^{\prime}\left\lvert R_{n}^{\prime}x^{\prime}\right\rvert\leq J^{\prime}R_{n}^{\prime}y^{\prime}$ in $X^{\prime}$ for each $n$.
Since the cone in $X$ is generating, the cone in $X^{\prime}$ is normal [1, Theorem 2.42] and hence, the previous inequality together with boundedness of the sequence $(J^{\prime}R_{n}^{\prime})$ gives the norm-boundedness of the sequence $\big{(}J^{\prime}\left\lvert R_{n}^{\prime}x^{\prime}\right\rvert\big{)}$ as well.
By the Banach-Alaoglu theorem, we can find a weak-convergent subsequence $\big{(}J^{\prime}\left\lvert R_{n_{j}}^{\prime}x^{\prime}\right\rvert\big{)}$ with limit $s^{\prime}$ in $X^{\prime}$. As the cone $X^{\prime}_{+}$ is weak-closed, it follows that $s^{\prime}\geq 0$. Let us show that $s^{\prime}$ is the supremum of ±$\pm x^{\prime}$ in $\operatorname{span}X^{\prime}_{+}$: for every upper bound $u^{\prime}$ of ±$\pm x^{\prime}$ in $\operatorname{span}X^{\prime}_{+}$ – repeating the above argument with $u^{\prime}$ instead of $y^{\prime}$ – we get
$J^{\prime}\left\lvert R_{n_{j}}^{\prime}x^{\prime}\right\rvert\leq J^{\prime}R_{n_{j}}^{\prime}u^{\prime}$
in $X^{\prime}$ for each index $j$. Employing weak-convergence of $(J^{\prime}R_{n}^{\prime})$ and weak-closedness of $X^{\prime}_{+}$, it thus follows that $s^{\prime}\leq u^{\prime}$, so $s^{\prime}$ is indeed the least upper bound of ±$\pm x^{\prime}$ in $\operatorname{span}X^{\prime}_{+}$.
Step 2: The span is a monotonically complete Banach lattice. As noted in Step 1, the cone in $X^{\prime}$ is normal. Hence, $X^{\prime}_{+}$ is also normal with respect to $\left\lVert\mathord{\,\cdot\,}\right\rVert_{\operatorname{span}X^{\prime}_{+}}$. Therefore, $\operatorname{span}X^{\prime}_{+}$ is an ordered Banach space with a normal cone and a vector lattice at the same time. For such spaces it is easy to check that they have an equivalent norm that turns them into a Banach lattice. To show that $\operatorname{span}X^{\prime}_{+}$ is monotonically complete, consider an increasing net $(x^{\prime}_{j})$ in $X^{\prime}_{+}$ that is norm-bounded in $\operatorname{span}X^{\prime}_{+}$. We need to show that it has a supremum in $\operatorname{span}X^{\prime}_{+}$. The net $(x^{\prime}_{j})$ is in particular norm-bounded in $X^{\prime}$ and hence, as it is increasing, it converges weak to a vector $x^{\prime}\in X^{\prime}_{+}$. It is easy to check that $x^{\prime}$ is the supremum of $(x^{\prime}_{j})$ in $X^{\prime}$ and hence in $\operatorname{span}X^{\prime}_{+}$.
Step 3: The span is a KB-space if $X$ is reflexive. If $X$ is reflexive, so is $X^{\prime}$. Moreover, as noted above the cone of $X^{\prime}_{+}$ is normal. Therefore, every increasing norm-bounded net in $X^{\prime}_{+}$ is norm-convergent in $X^{\prime}$. [1, Theorem 2.45]. Applying Lemma 2.2, we immediately obtain that every increasing norm-bounded net in $X^{\prime}_{+}$ is also norm-convergent with respect to $\left\lVert\mathord{\,\cdot\,}\right\rVert_{\operatorname{span}X^{\prime}_{+}}$. In other words, $\operatorname{span}X^{\prime}_{+}$ is a KB-space. ∎
Remark 2.3.
From Step 3 in the proof of Theorem 2.1, we see that the assumption $X$ is reflexive in Theorem 2.1 can be replaced with “every increasing norm-bounded net in $X^{\prime}_{+}$ is norm-convergent”.
This assumption is, for instance, satisfied if $X_{+}$ has non-empty interior, since then there exists an equivalent norm on $X^{\prime}$ that is additive on the positive cone, which in turn gives that every increasing norm-bounded net in $X^{\prime}_{+}$ is a Cauchy net.
Let us also explicitly state a version of Theorem 2.1 where the conclusion refers to the spaces themselves rather than to their duals. Due to the dualization, the roles of $X$ and $Z$ in the assumptions of the following corollary are swapped compared to Theorem 2.1.
Corollary 2.4.
Let $X$ be a Banach lattice and let $Z$ be a reflexive ordered Banach space with a normal cone. Assume that there exists an operator $J\in\mathcal{L}(X,Z)_{+}$ and a sequence $(R_{n})\subseteq\mathcal{L}(Z,X)_{+}$ such that $(JR_{n})\subseteq\mathcal{L}(Z)$ converges to the $\operatorname{id}_{Z}$ with respect to the weak operator topology. Then $\operatorname{span}(Z_{+})=Z_{+}-Z_{+}$ is a KB-space with respect to a norm equivalent to $\left\lVert\mathord{\,\cdot\,}\right\rVert_{\operatorname{span}Z_{+}}$.
Proof.
The dual space $\tilde{Z}:=X^{\prime}$ is a Banach lattice and the cone in $\tilde{X}:=Z^{\prime}$ is generating since the cone in $Z$ is normal [1, Theorem 2.40].
Moreover, the sequence $(R_{n}^{\prime}J^{\prime})\subseteq\mathcal{L}(Z^{\prime})$ converges to $\operatorname{id}_{Z^{\prime}}$ with respect to the weak-operator topology. Since $Z^{\prime}$, is reflexive, this topology coincides with the weak operator topology. So we can apply Theorem 2.1 to the spaces $\tilde{X}$ and $\tilde{Z}$ and the operators $J^{\prime}$ and $R_{n}^{\prime}$ to conclude that $\operatorname{span}\tilde{X}^{\prime}_{+}=\operatorname{span}Z_{+}$ is a KB-space. ∎
Even though surprising, $\operatorname{span}Z_{+}$ need not be reflexive in Corollary 2.4; see Example 4.2. This also shows that $\operatorname{span}X^{\prime}_{+}$ need not be reflexive in Theorem 2.1 even if $X$ is so.
A Banach lattice $X$ is said to have order continuous norm if every increasing and order-bounded net (equivalently, sequence) in $X_{+}$ is norm-convergent. Classical examples are $L^{p}$-spaces for $p\in[1,\infty)$ as well as the space $c_{0}$. KB-spaces have order continuous norm, but the converse is not true as the example $c_{0}$ shows. We refer to [14] for extensive literature on Banach lattices with order continuous norms.
Theorem 2.5.
Let $X$ be a Banach lattice with order continuous norm and let $Z$ be an ordered Banach space. Assume that there exists a positive bijection $S\in\mathcal{L}(Z,X)$, a positive operator $J\in\mathcal{L}(X,Z)$ and a sequence of positive operators $(R_{n})\subseteq\mathcal{L}(Z,X)$ such that $(JR_{n})\subseteq\mathcal{L}(Z)$ converges to the $\operatorname{id}_{Z}$ with respect to the strong operator topology.
Then $\operatorname{span}(Z_{+})$ is a Banach lattice with order continuous norm with respect to a norm that is equivalent to $\left\lVert\mathord{\,\cdot\,}\right\rVert_{\operatorname{span}Z_{+}}$. If $X$ is a KB-space, then so is $\operatorname{span}(Z_{+})$.
Note that we do not assume $S$ to be bipositive – i.e., the inverse operator $S^{-1}\in\mathcal{L}(X,Z)$ need not be positive. This will be important for our application to extrapolation spaces in Section 4.
We point out that Theorem 2.5 gives that $\operatorname{span}Z_{+}$ is a Banach lattice if $X$ is a Banach lattice with order continuous norm, while Corollary 2.4 requires $X$ to be reflexive to get the same conclusion. Similarly, Theorem 2.5 gives that $\operatorname{span}Z_{+}$ is a KB-space Banach lattice if $X$ is so, while Corollary 2.4 again requires $X$ to be reflexive to get the same conclusion. The prize that one pays for this weaker assumption in Theorem 2.5 is that one also needs the existence of the bijective positive operator $S$ and that the sequence $(JR_{n})$ is required to converge with respect to the strong instead of the weak operator topology.
Proof of Theorem 2.5.
Step 1: We first make the following auxiliary observation: If $(f_{n})$ and $(g_{n})$ are two sequences in $Z$ such that $0\leq f_{n}\leq g_{n}$ for all indices $n$ and if $(g_{n})$ converges with respect to the norm in $Z$ to a vector $g\in Z$, then $(f_{n})$ has a subsequence that converges weakly in $Z$.
Indeed, after replacing sequences with subsequences, we may assume that one has $\sum_{n=1}^{\infty}\left\lVert g_{n}-g\right\rVert_{Z}<\infty$. Then the values $\left\lVert Sg_{n}-Sg\right\rVert_{X}$ are also summable, so the series $h:=\sum_{n=1}^{\infty}\left\lvert Sg_{n}-Sg\right\rvert$ converges absolutely in $X$. One obtains that $0\leq Sf_{n}\leq Sg_{n}\leq h+Sg$ for all indices $n$, where the first two inequalities follow from the positivity of $S$ and the last inequality from the definition of $h$. So the sequence $(Sf_{n})$ is contained in an order interval of $X$ and since $X$ has order continuous norm, all order intervals in $X$ are weakly compact [14, Theorem 1.12]. Thus $(Sf_{n})$ has a subsequence that converges weakly in $X$. Since the inverse operator $S^{-1}:X\to Z$ (which might not be positive) is continuous by the bounded inverse theorem and is thus weakly continuous from $X$ to $Z$, it follows that $(f_{n})$ does indeed have a subsequence that converges weakly in $Z$.
Step 2: The span is a lattice. To see this, let $z\in\operatorname{span}Z_{+}$. It suffices to show that ±$\pm z$ has a supremum in $\operatorname{span}Z_{+}$. As $z\in Z_{+}-Z_{+}$, there exists $y\in Z_{+}$ such that ±$\pm z\leq y$. Then we have ±$\pm R_{n}z\leq R_{n}y$ and hence $0\leq J\left\lvert R_{n}z\right\rvert\leq JR_{n}y$ for all indices $n$. By assumption, $(JR_{n}y)$ converges to $y$ with respect to the norm in $Z$, and thus the first part of the proof shows that $(J\left\lvert R_{n}z\right\rvert)$ has a subsequence $(J\left\lvert R_{n_{j}}z\right\rvert)$ that converges weakly to an element $s\in Z$. As $Z_{+}$ is closed and convex it is weakly closed, so $s\in Z_{+}$. We show that $s$ is the supremum of ±$\pm z$ within $\operatorname{span}(Z_{+})$. For every upper bound $u\in\operatorname{span}(Z_{+})$ of ±$\pm z$ – repeating the above argument with $u$ instead of $y$ – we get
$J\left\lvert R_{n_{j}}z\right\rvert\leq JR_{n_{j}}u$
for all indices $j$. Taking weak limits in $Z$ and using again that the cone $Z_{+}$ is weakly closed in $Z$ we get $s\leq u$. Hence, $s$ is indeed the least upper bound of ±$\pm z$ within $\operatorname{span}(Z_{+})$.
Step 3: The span is a Banach lattice with order continuous norm. We first show that the cone $Z_{+}$ in $Z$ is normal. To see this, let $y,z\in Z$ satisfy $0\leq y\leq z$. Then we have $0\leq Sy\leq Sz$ in $X$ and hence, as $X$ is a Banach lattice, $\left\lVert Sy\right\rVert\leq\left\lVert Sz\right\rVert$. The inverse operator $S^{-1}$ is bounded by the bounded inverse theorem and thus,
$\left\lVert y\right\rVert\leq\left\lVert S^{-1}\right\rVert\left\lVert Sy\right\rVert\leq\left\lVert S^{-1}\right\rVert\left\lVert Sz\right\rVert\leq\left\lVert S^{-1}\right\rVert\left\lVert S\right\rVert\left\lVert z\right\rVert.$
Therefore, $Z_{+}$ is normal in $Z$, and in turn, also in $\operatorname{span}Z_{+}$; see the discussion after (1.1). So $\operatorname{span}(Z_{+})$ is an ordered Banach space with a normal cone and, at the same time, a vector lattice. Hence, it is a Banach lattice with respect to an equivalent norm.
In order to show that the norm in $\operatorname{span}Z_{+}$ is order continuous, let $(z_{j})$ be an increasing order-bounded net in $Z_{+}$; we need to show that it is norm-convergent in $\operatorname{span}Z_{+}$. The net $(Sz_{j})$ in $X_{+}$ is also increasing and order-bounded, so it is norm-convergent in $X$, since $X$ has order continuous norm. By the bounded inverse theorem, we obtain that the net $(z_{j})$ is norm-convergent with respect to the norm in $Z$. As $(z_{j})$ was increasing, we may apply Lemma 2.2 to assert the claim.
Step 4: The span is a KB-space if $X$ is so. Since every increasing norm-bounded net in $\operatorname{span}(Z_{+})$ is also increasing and norm-bounded in $Z$, this is yet another application of Lemma 2.2. ∎
3. Sobolev spaces of negative order
Let $p,q\in(1,\infty)$ be Hölder conjugates and let $d,k\geq 1$ be integers. We endow the Sobolev space $W^{k,p}(\mathbb{R}^{d})$ with the usual order inherited from $L^{p}(\mathbb{R}^{d})$. This is an ordered Banach space with generating cone (even for $p\in[1,\infty]$, see the introduction of [13]); the cone is not normal as can easily be seen by considering rapidly oscillating functions. Furthermore, $W^{k,p}(\mathbb{R}^{d})$ is a vector lattice if and only if $k=1$, see, for instance, [9, Lemma 7.6] and [2, Example 2.3(d)]; however, even for the case $k=1$ the space is not a Banach lattice due to the cone not being normal.
Let us now use Theorem 2.1 to demonstrate that the situation is different for Sobolev spaces of negative order. The space $W^{-k,p}(\mathbb{R}^{d})$ is the dual space of $W^{k,q}(\mathbb{R}^{d})$ and it thus contains $L^{p}(\mathbb{R}^{d})$ in a canonical way. We endow $W^{-k,p}(\mathbb{R}^{d})$ with the cone $W^{-k,p}(\mathbb{R}^{d})_{+}$ that is defined to be the dual cone of $W^{k,q}(\mathbb{R}^{d})_{+}$; this dual cone can easily be seen to coincide with the closure of $L^{p}(\mathbb{R}^{d})_{+}$ in $W^{-k,p}(\mathbb{R}^{d})$. As noted above, the cone $W^{k,q}(\mathbb{R}^{d})_{+}$ is not normal. Hence, by duality the cone $W^{-k,p}(\mathbb{R}^{d})_{+}$ is not generating in $W^{-k,p}(\mathbb{R}^{d})$ [1, Theorem 2.40]. So in particular $W^{-k,p}(\mathbb{R}^{d})$ cannot be vector lattice. Nevertheless, we show that the span of the cone in this space is still a vector lattice – and even a Banach lattice when endowed with an appropriate norm.
Theorem 3.1.
Let $p\in(1,\infty)$ and let $d,k\geq 1$ be integers. The span of $W^{-k,p}(\mathbb{R}^{d})_{+}$ in $W^{-k,p}(\mathbb{R}^{d})$ is a KB-space (in particular, a Banach lattice) with respect to a norm equivalent to $\left\lVert\mathord{\,\cdot\,}\right\rVert_{\operatorname{span}W^{-k,p}(\mathbb{R}^{d})_{+}}$.
Proof.
Let $q\in(1,\infty)$ be the Hölder conjugate of $p$ and let $J:X:=W^{k,q}(\mathbb{R}^{d})\to L^{q}(\mathbb{R}^{d})=:Z$ be the canonical embedding. Let $\rho:\mathbb{R}^{d}\to[0,\infty)$ be a test function with integral $1$ and define the mollifier sequence $(\rho_{n})$ by assumption $\rho_{n}:=n\rho(n\mathord{\,\cdot\,})$ for each integer $n\geq 1$. For each $n$, let $R_{n}:Z\to X$ be given by $R_{n}f=\rho_{n}\star f$ for every $f\in Z$. Then $\big{(}R_{n}J\big{)}$ converges strongly to the identity on $X$. Hence, the claim follows by an application of Theorem 2.1 since $W^{-k,p}(\mathbb{R}^{d})=X^{\prime}$. ∎
4. Extrapolation spaces of positive semigroups
Let $X$ be a Banach lattice and let $(T(t))_{t\geq 0}$ be a $C_{0}$-semigroup on $X$ that is positive, which means that each operator $T(t)$ is positive. Let $A$ denote the generator of $(T(t))_{t\geq 0}$; which is a closed and densely defined operator. For the general theory of $C_{0}$-semigroups, we refer to [6] and for positive $C_{0}$-semigroups specifically to [12, 5].
Fix a real number $\lambda$ that is larger than the real part of every spectral value of $A$. The extrapolation space $X_{-1}$ is defined as follows: for every $x\in X$ one sets $\left\lVert x\right\rVert_{-1}:=\left\lVert(\lambda-A)^{-1}x\right\rVert_{X}$ and then defines $X_{-1}$ to be the completion of $X$ with respect to this norm. Note that $X_{-1}$ does not depend on the choice of $\lambda$ since all norms that one obtains in this way are equivalent. One can extend the semigroup generator $A$ to an operator $A_{-1}:X\subseteq X_{-1}\to X_{-1}$. The resolvent can also be extended to an operator from $X_{-1}$ to $X$ that then intertwines $A$ and $A_{-1}$.
The space $X_{-1}$ plays, for instance, a role in perturbation and linear systems theory. In order to study positive perturbations the authors of [4] introduced a cone $X_{-1,+}$ in $X_{-1}$ which they defined to be the closure of $X_{+}$ within $X_{-1}$. This is indeed a cone and satisfies $X_{-1,+}\cap X=X_{+}$ [4, Proposition 2.3]. The extended resolvent operators $(\mu-A_{-1})^{-1}$ are positive from $X_{-1}$ to $X$ for each $\mu$ that is larger than the real part of all spectral values of $A$ [4, Remark 2.2]. Various order properties of the cone $X_{-1,+}$, even in the case that $X$ is an ordered Banach space were recently proved in [3, Section 2.2].
As mentioned in the introduction, $X_{-1,+}$ is usually not generating in $X_{-1}$ and so $X_{-1}$ is not a vector lattice. However, because every cone of an ordered Banach space is generating in its span, it is reasonable to ask if the span of $X_{-1,+}$ is a vector lattice. This is indeed often the case, as the following consequence of Theorem 2.5 shows.
Theorem 4.1.
Let $(T(t))_{t\geq 0}$ be a positive $C_{0}$-semigroup on a Banach lattice $X$ with order continuous norm. Then $\operatorname{span}(X_{-1,+})$ is – with respect to a norm equivalent to $\left\lVert\mathord{\,\cdot\,}\right\rVert_{\operatorname{span}X_{-1,+}}$ – a Banach lattice with order continuous norm. If $X$ is a KB-space, then so is $\operatorname{span}(X_{-1,+})$.
Proof.
Apply Theorem 2.5 to the space $Z:=X_{-1}$ with $J:X\to Z$ being the canonical embedding, $R_{n}:=n(n-A_{-1})^{-1}$ for all sufficiently large integers $n$, and $S:=R_{n_{0}}$ for some fixed $n_{0}$. ∎
In some situations one can concretely determine the space $\operatorname{span}X_{-1,+}$:
Example 4.2.
Let $\Omega\subseteq\mathbb{R}^{d}$ be a bounded domain with Lipschitz boundary, let $p,q\in(1,\infty)$ be Hölder conjugates, and let $\Delta$ denote the Neumann Laplace operator on $X=L^{p}(\Omega)$. If $d<2q$, then $\operatorname{span}X_{-1,+}$ is precisely the space of finite Borel measures on $\overline{\Omega}$; see [3, Example 2.13(b)]. Note that, since positive operators on ordered Banach spaces with generating cone are automatically continuous [1, Theorem 2.32], all complete norms that turn $\operatorname{span}X_{-1,+}$ into an ordered Banach space, are equivalent. Hence, the norm $\left\lVert\mathord{\,\cdot\,}\right\rVert_{\operatorname{span}X_{-1,+}}$ is equivalent to the total variation norm in this example and consequently, the space $\operatorname{span}X_{-1,+}$ is not reflexive despite $X$ and $Z$ both being reflexive.
Since the assumptions of Corollary 2.4 are also satisfied in this example (with $Z:=X_{-1}$) this also shows that the conclusion of Corollary 2.4 cannot be strengthened to give that $\operatorname{span}Z_{+}$ is reflexive.
5. Outlook
Assume that we are in the situation of Theorem 2.5, but that the Banach lattice $X$ is not required to have order continuous norm. We do not know whether $\operatorname{span}Z_{+}$ still needs to be a Banach lattice in this case (the arguments in the proof of Theorem 2.5 show that it suffices to prove that $\operatorname{span}Z_{+}$ is a vector lattice). This is the first of the following two open problems.
Open Problems 5.1.
[…]
    Let $X$ be a Banach lattice and let $Z$ be an ordered Banach space. Assume that there exists a positive bijection $S\in\mathcal{L}(Z,X)$, a positive operator $J\in\mathcal{L}(X,Z)$ and a sequence of positive operators $(R_{n})\subseteq\mathcal{L}(Z,X)$ such that $(JR_{n})\subseteq\mathcal{L}(Z)$ converges to the $\operatorname{id}_{Z}$ with respect to the strong operator topology.
    Does it follow that $\operatorname{span}Z_{+}$ is a vector lattice (and thus a Banach lattice with respect to a norm equivalent to $\left\lVert\mathord{\,\cdot\,}\right\rVert_{\operatorname{span}Z_{+}})$?
[…]
    If the answer to problem (a) is negative, does it become positive if the mapping $J$ is compact?
As explained before the problem statement, Theorem 2.5 says that the answer to problem (a) is positive if $X$ has order continuous norm. Two concrete examples in which $X$ does not have order continuous norm are given in [4, Examples 5.1 and 5.3]. In both examples, $X$ is a space of continuous functions on an interval (so it does not have order continuous norm) and $Z:=X_{-1}$ is the extrapolation space of a positive $C_{0}$-semigroup on $X$. However, in both examples the positive cone is computed explicitly in [4, formulas (5.4) and (5.6)] and those results show that the span of the cone is indeed a lattice. In both examples the canonical embedding from $X$ into $X_{-1}$ is compact (since the generator of the semigroup has compact resolvent in both cases). This might be an indication that the answer to problem (b) above is positive.
Acknowledgements
The first author was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 523942381. The second author is indebted to the Department of Applied Mathematics of the University of Twente for a very pleasant stay during which some parts of the work on this article were done. The authors are grateful to Lassi Paunonen for useful comments on the manuscript.